OpenStudy (anonymous):

Simplyfying radicals anyone? They're really damn hard. I can post paper if anyone's up to it! Will give medal and fan!!! Please help,

OpenStudy (_xoxo_hearts_):

sure whats your question ?

OpenStudy (anonymous):

I need to get a 3.5 in this to graduate high school, and I've missed a couple days due to family issues.. So I'm screwed. :/ I'd super appreciate the help

OpenStudy (_xoxo_hearts_):

okay I understand ^.^ same actually

OpenStudy (anonymous):

yeah, thank you! This stuff is just really hard when you don't know what you're doing

OpenStudy (_xoxo_hearts_):

okay what number are you working on ?

OpenStudy (_xoxo_hearts_):

or is it all of them?

OpenStudy (anonymous):

All of them..

OpenStudy (_xoxo_hearts_):

okay do you want the answers or do you want to work them

OpenStudy (anonymous):

Answers would be the easiest, so then I could just answer them! Thank you!

OpenStudy (_xoxo_hearts_):

sure im dm them to you

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Simplifying a radical is all about taking out number from inside it to outside it, so that it multiplies the radical. First, take account of the index of the radical and make each and every term inside the radical as a factor. For instance, if you have sqrt75, it simplifies to 5sqrt3, because the factore of 75 are 5*5*3. Then you pile the to 5^2*3. Having that 5 has a power of 2 and so the index, then they can be simplified, because 2/2=1, and this one, will be the power of the 5 as it will move out of the radical. For numbers that do not have divisible numbers in relation to the index of the radical shall remain inside it. That is why the 3, which has power of 1 does not simplify with 2, hence stays inside. the reason why sqrt75= 5sqrt3. Using one of your given exercises: sqrt20m^5n= sqrt2^2*5*m^4*m*n= 2*m^2sqrt5mn

OpenStudy (anonymous):

yeah, it's just super difficult, thanks tho man

OpenStudy (anonymous):

Just take step by step profundly and any question, ask. I am glad to help anyway.

OpenStudy (anonymous):

okay, thank you!

OpenStudy (anonymous):

Someone help, I have no idea how todo this